# Connectedness of the Gromov boundary of fine curve graphs

### v2

@@ 5 Group action on the boundary
The action of $\operatorname{Homeo}(S)$ on $\partial\mathcal{C}^{\dagger}(S)$ is continuous.
@@
It suffices to show that whenever given a sequence $g_{n}\to\operatorname{Id}$ as $n\to\infty$, for any $M>0$ and $\xi\in\partial\mathcal{C}^{\dagger}(S)$, there exists $M^{\prime}>0$ and $N>0$ such that if $\langle\xi,\eta\rangle_{o}>M^{\prime}$, then $\langle\xi,g_{n}\eta\rangle_{o}>M$ for any $n>N$. Indeed, since the open sets on the boundary is also generated by shadows defined in (4), we can find an $\alpha\in\mathcal{C}^{\dagger}(S)$ such that whenever $\eta\in\partial\mathcal{C}^{\dagger}(S)$ satisfies $\langle\alpha,\eta\rangle_{o}>R$, we will have $\langle\eta,\xi\rangle_{o}>M$. Moreover, we can assume that $\langle\alpha,\xi\rangle_{o}$ is much larger than $R$. Since the topology on $\operatorname{Homeo}(S)$ is the compact-open topology, there exists $N>0$ such that $g_{n}\alpha$ is contained in a uniformly small neighbourhood of $\alpha$ for every $n>N$. This implies that for any curve $\gamma\subset S$ that is disjoint from this small neighbourhood and a fortiori from $\alpha$, $\gamma$ is also disjoint from $g_{n}\alpha$ for all $n>N$. Hence $d^{\dagger}(\alpha,g_{n}\alpha)\leq 2$ for all $n>N$. The same arguments hold for $o$ so that $d^{\dagger}(o,g_{n}o)\leq 2$ for $n>N$. Let $\eta\in\partial\mathcal{C}^{\dagger}(S)$ such that $\langle\eta,\xi\rangle_{o}>M^{\prime}$. We compute by (2) and (3)
@@
In view of Proposition 5.3 and Lemma 5.2, it suffices to show that the action of $\operatorname{Homeo}(S)$ on $\mathcal{C}^{\dagger}(S)$ is of general type. To prove so, it suffices to fix a hyperbolic metric on $S$ and take two independent pseudo-Anosov isometric homeomorphisms on $S$, which will yield two independent hyperbolic isometries on $\mathcal{C}^{\dagger}(S)$ as their action of a subgraph of $\mathcal{C}^{\dagger}(S)$ isometric to $\mathcal{C}(S)$ is of general type (see [MM99]). ∎
@@
To conclude this section, we wish to point out in the interests of curiosity that if $S$ is a compact surface, then $\operatorname{Homeo}(S)$ is a Polish group, i.e. a completely metrizable and separable topological group, and its continuous minimal actions on compacta are of special interests. For example, [GTZ21] shows that its universal minimal flow, a compact space with a minimal $\operatorname{Homeo}(S)$-action that admits any compact minimal space for $\operatorname{Homeo}(S)$-actions as a factor, is not metrizable.

### v4

@@ 5 Connectivity properties
5.B. Connectedness of the spheres.
In this subsection, we address the connectedness properties of spheres within the fine curve graph, building on techniques developed in [wright2023spheres]:
Lemma 5.8 (Lemma 2.1, [wright2023spheres]).
Let $\Gamma$ be an arbitrary graph and $o\in\Gamma$ be arbitrary. Fix $w>0$. Let $S_{r}$ be the $r$-sphere in $\Gamma$. Suppose in addition that for every $r>0$, the following conditions hold:
  1. (G2’)
    For every $z\in S_{r}$ and $x,y\in S_{r+1}\cap B_{z}(1)$ there exists a path
    $x=x_{0},\ x_{1},\ \dots,\ x_{\ell}=y$
    with
    $x_{i}\in S_{r+1}\cup\dots\cup S_{r+w}$
    for $0\leq i\leq\ell$.
  2. (G3’)
    For every adjacent pair $x,y\in S_{r}$ there exists a path
    $x=x_{0},\ x_{1},\ \dots,\ x_{\ell}=y$
    with
    $x_{i}\in S_{r+1}\cup\dots\cup S_{r+w}$
    for $0<i<\ell$. Then for any $r>0$, the union $S_{r}\cup\dots\cup S_{r+w-1}$ is connected.
Remark 5.9.
We remark that the condition (G2) (respectively (G3)) is stronger than the condition (G2’) (respectively (G3’)).
In order to show the connectedness of the spheres in $\mathcal{C}^{\dagger}(S)$, we will first check the connectedness of the spheres in $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$.
Lemma 5.10.
Let $S$ be a surface as given above. For any positive integer $r>0$, the sphere $\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ is connected.
Proof.
Lemma 5.5 and Lemma 5.6 show that the graph $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$ verifies (G2) and (G3), a fortiori (G2’) and (G3’). By Lemma 5.8, the sphere $\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ is connected. ∎
First, we observe that, by Lemma 3.12,
$\operatorname{\mathcal{S}_{o,\pitchfork}}(r)\subset\mathcal{S}_{o}(r)=\{\alpha\in\operatorname{\mathcal{NC}}^{\dagger}(S)\colon d^{\dagger}(\alpha,o)=r\}.$ (5.2)
Hence the sphere $\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ forms a subgraph in the sphere $\mathcal{S}_{o}(r)$. Although Proposition 3.11 already shows that $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$ is $1$-dense in $\operatorname{\mathcal{NC}}^{\dagger}(S)$, the following lemma further demonstrates that it is in fact $1$-dense at each $r$-sphere.
Lemma 5.11.
Let $S$ be a surface as given above. For any $\alpha\in\mathcal{S}_{o}(r)$, there exists $\alpha^{\prime}\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ such that $d^{\dagger}(\alpha,\alpha^{\prime})=1$, i.e., $\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ is 1-dense in $\mathcal{S}_{o}(r)$.
Proof.
Let $o,v_{1},v_{2},\dots,v_{r-1},\alpha$ be a geodesic path in $\operatorname{\mathcal{NC}}^{\dagger}(S)$ connecting $o$ to $\alpha$. There is a finite subset $P\subset S$ and $\alpha_{1}\in\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$ both disjoint from $\alpha$ such that $\alpha_{1}$ is also disjoint from $v_{r-1}$, isotopic to $\alpha$ on $S-P$ and in minimal position with $o$ on $S-P$. By Lemma 3.12 and triangle inequality,
$r-1=d_{o}^{\dagger}(o,\alpha)-d_{o}^{\dagger}(\alpha,\alpha_{1})\leq d^{\dagger}(o,\alpha_{1})=d_{o}^{\dagger}(o,\alpha_{1})\leq d^{\dagger}_{o}(o,v_{r-1})+d^{\dagger}_{o}(v_{r-1},\alpha_{1})=r.$
If $d_{o}^{\dagger}(o,\alpha_{1})=r$, then there is nothing to prove. So we assume that $d_{o}^{\dagger}(o,\alpha_{1})=r-1$.
By Lemma 5.6, we can take a $\beta\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ disjoint from both $v_{r-1}$ and $\alpha_{1}$. Note that $\alpha\cap o$ is not discrete. Since $\alpha$ and $o$ are proper images, we can take a family $\big{(}U_{i}\big{)}_{i=1,\dots,m}$ of small open subsets on $S$ such that $U_{i}\cap U_{j}=\emptyset$ for $i\neq j$, $U\coloneqq\bigcup_{i=1}^{m}U_{i}$ does not contain the entire $o$ and $(\alpha\cap o)-U$ is discrete, i.e., $\alpha$ and $o$ are transverse outside of $U$. Hence we can similarly isotope $\beta$ as in Lemma 3.9 so that $\beta$ is transverse to both $o$ and $\alpha$ outside of $U$. Finally, by isotoping $\beta$ away from $U$ and preserving the intersection pattern, we can obtain a $\beta$ that is transverse to both $\alpha$ and $o$. Note that if $U$ is small enough, then $\beta$ can remain in minimal position with $o$ on $S-P$, guaranteeing that $d^{\dagger}_{o}(o,\beta)=r$ after a quick application of Lemma 3.12 and Proposition 3.7.
If $\beta$ is already disjoint from $\alpha$, then $\beta$ will be the $\alpha^{\prime}$ we are searching for. Otherwise, there exists at least one bigon bounded by subarcs of $\alpha$ and $\beta$ on $S-P$. By isotope the subarc of $\beta$ along the bigon to the other side of $\alpha$ and then perturb the curve, we can obtain a curve $\beta_{1}$ with strictly fewer intersection points with $\alpha$, disjoint from $\beta$ and remaining transverse to $o$. Since this bigon does not contain any points from $P$, $\beta_{1}$ stays in minimal position with $o$. Hence by Lemma 3.12 and Proposition 3.7 again, $\beta_{1}\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$. Because $\alpha_{1}\in[\alpha]_{S-P}$ and $\beta\in[\beta]_{S-P}$ are disjoint, we can inductively construct a path $\beta_{1},\beta_{2},\cdots,\beta_{n}\subset\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ until $\beta_{n}$ and $\alpha$ are in minimal position (i.e. no bigon), which in this case will be disjoint. Thus $\alpha^{\prime}=\beta_{n}$ is the curve we are looking for.
Finally, we remark that all the perturbation and isotopy can be done on the subsurface $S-\gamma$ if $\alpha$ is disjoint from $\gamma$. This means that $\alpha^{\prime}$ can be taken disjoint from $\gamma$. ∎
Note that the perturbation in the proof of Lemma 5.11 can also be carried out for separating curves. Using the $d_{\mathcal{C}}^{s}(S-P)$-distance estimation for curve in minimal position instead of $d_{\mathcal{NC}(S-P)}$, we can conclude the same result for separating curves:
Lemma 5.12.
Let $S$ be a surface as given above. For any $\alpha\in\mathcal{C}^{\dagger}(S)$ with $d^{\dagger}(o,\alpha)=r>0$, there exists $\alpha^{\prime}\in\mathcal{C}^{\dagger}(S)$ transverse to $o$ such that $d^{\dagger}(\alpha,\alpha^{\prime})=1$ and $d^{\dagger}(o,\alpha^{\prime})=r$.
Now we are able to conclude Theorem 1.2.
Proof of Theorem 1.2.
Let $r>0$ be any positive integer. By Lemma 5.10, the sphere $\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ in $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$ is connected. As $\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ is $1$-dense in the sphere $\mathcal{S}_{o}(r)$ in $\mathcal{NC}^{\dagger}(S)$ by Lemma 5.11, $\mathcal{S}_{o}(r)$ is also connected. ∎
As an corollary, we can deduce a connectedness of thickened spheres in the fine curve graph as below:
Corollary 5.13.
Let $S$ be as given above. Then for any integer $r>0$ and any $o\in\mathcal{C}^{\dagger}(S)$, then the thickened sphere
$\Sigma_{r}\coloneqq\left\{\gamma\in\mathcal{C}^{\dagger}(S)\colon d^{\dagger}(o,\gamma)=r,r+1\right\}$
is connected.
Proof.
Let $r>0$ be any positive integer. Let us consider the subgraph
$\operatorname{\mathcal{C}^{\dagger}_{o,\pitchfork}}(S)\coloneqq\left\{\alpha\in\mathcal{C}^{\dagger}(S)\colon\alpha\text{ is transverse to }o\right\}\subset\mathcal{C}^{\dagger}(S)$
of curves transverse to $o$. By Remark 3.13, the thickened sphere in $\operatorname{\mathcal{C}^{\dagger}_{o,\pitchfork}}(S)$
$\Sigma^{\pitchfork}_{r}\coloneqq\left\{\gamma\in\operatorname{\mathcal{C}^{\dagger}_{o,\pitchfork}}(S)\colon d^{\dagger}(o,\gamma)=r,r+1\right\}$
is a subgraph of $\Sigma_{r}$. By the virtue of Lemma 5.12, it suffices to verify that $\Sigma^{\pitchfork}_{r}$ is connected.
Indeed, let $\alpha,\beta\in\operatorname{\mathcal{C}^{\dagger}_{o,\pitchfork}}(S)$ be any disjoint pair with $d^{\dagger}(o,\alpha)=d^{\dagger}(o,\beta)=r$. Similar to the previous cases, we can find a non-separating curve $\alpha^{\prime}\in\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$ disjoint from $\alpha$ such that $d_{Y}\big{(}\pi_{Y}(o),\pi_{Y}(\alpha^{\prime})\big{)}>M$, where $M$ is the constant from Theorem 4.12 and $Y=S-\alpha$. This will make $d^{\dagger}(o,\alpha^{\prime})=r+1$, or equivalently $\alpha^{\prime}\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r+1)$. Similarly, we can also find a $\beta^{\prime}\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r+1)$ disjoint from $\beta$. Appealing to Lemma 5.10, we can find a path on $\operatorname{\mathcal{S}_{o,\pitchfork}}(r+1)$ connecting $\alpha^{\prime}$ to $\beta^{\prime}$. Now the concatenating path
$\alpha\text{ --- }\alpha^{\prime}\rightsquigarrow\beta^{\prime}\text{ --- }\beta$
verifies the condition (G3’) for $\operatorname{\mathcal{C}^{\dagger}_{o,\pitchfork}}(S)$ with $w=1$. The same arguments can also allow us to verify (G2’) for $\operatorname{\mathcal{C}^{\dagger}_{o,\pitchfork}}(S)$ with $w=2$. Hence Lemma 5.8 applies to complete the proof. ∎
@@ 6 Group action on the boundary
The action of $\operatorname{Homeo}(S)$ on $\partial\mathcal{C}^{\dagger}(S)$ is (uniformly) continuous.
@@
It suffices to show that whenever given a sequence $g_{n}\to\operatorname{Id}$ as $n\to\infty$, for any $M>0$ and $\xi\in\partial\mathcal{C}^{\dagger}(S)$, there exists $M^{\prime}>0$ and $N>0$ such that if $\langle\xi,\eta\rangle_{o}>M^{\prime}$, then $\langle\xi,g_{n}\eta\rangle_{o}>M$ for any $n>N$.
Indeed, since the open sets on the boundary is also generated by shadows defined in (2.4), we can find an $\alpha\in\mathcal{C}^{\dagger}(S)$ such that whenever $\eta\in\partial\mathcal{C}^{\dagger}(S)$ satisfies $\langle\alpha,\eta\rangle_{o}>R$, we will have $\langle\eta,\xi\rangle_{o}>M$. Moreover, we can assume that $\langle\alpha,\xi\rangle_{o}$ is much larger than $R$.
Since the topology on $\operatorname{Homeo}(S)$ is the compact-open topology, there exists $N>0$ such that $g_{n}\alpha$ is contained in a uniformly small neighbourhood of $\alpha$ for every $n>N$. This implies that for any curve $\gamma\subset S$ that is disjoint from this small neighbourhood and a fortiori from $\alpha$, $\gamma$ is also disjoint from $g_{n}\alpha$ for all $n>N$. Hence $d^{\dagger}(\alpha,g_{n}\alpha)\leq 2$ for all $n>N$. The same arguments hold for $o$ so that $d^{\dagger}(o,g_{n}o)\leq 2$ for $n>N$. Let $\eta\in\partial\mathcal{C}^{\dagger}(S)$ such that $\langle\eta,\xi\rangle_{o}>M^{\prime}$. We compute by (2.2) and (2.3)
@@
In view of Proposition 6.3 and Lemma 6.2, it suffices to show that the action of $\operatorname{Homeo}(S)$ on $\mathcal{C}^{\dagger}(S)$ is of general type. To prove so, it suffices to take two independent pseudo-Anosov homeomorphisms on $S$, which will yield two independent hyperbolic isometries on $\mathcal{C}^{\dagger}(S)$ as their action of a subgraph of $\mathcal{C}^{\dagger}(S)$ isometric to $\mathcal{C}(S)$ is of general type (see [masur1999geometry]). ∎
@@
To conclude this section, we wish to point out in the interests of curiosity that $\operatorname{Homeo}(S)$ is a Polish group, i.e. a completely metrizable and separable topological group, and its continuous minimal actions on compacta are of special interests. For example, [gutman2021universal] shows that its universal minimal flow, a compact space with a minimal $\operatorname{Homeo}(S)$-action that admits any compact minimal space for $\operatorname{Homeo}(S)$-actions as a factor, is not metrizable.
[…]
Yusen Long
Université Paris-Saclay, Institut Mathématique d’Orsay, 91405 Orsay, France
Université Paris-Est Créteil, CNRS, LAMA UMR8050, F-94010 Creteil, France
Dong Tan
Guangxi University, Guangxi Center for Mathematical Research, 530004 Nanning, China
Email address: duzuizhe2013@foxmail.com